\begin{thebibliography}{99}
\bibitem{nbs}M. Abramowitz and I. A. Stegun,
{Handbook of Mathematical Functions, Washington, National Bureau of Standards
(1964).}
%\bibitem{apostol}T. M. Apostol,
%{Introduction to Analytic Number Theory, Springer Verlag, New York (1976);
%corrected fourth printing, 1995.}
%\bibitem{balazard}M. Balazard, E. Saias, and M. Yor,
%{Notes sur la fonction $\zeta$ de Riemann, 2, Advances in Mathematics 
%{\bf 143}, 284-287 (1999).}
\bibitem{berndt}B. C. Berndt,
{On the Hurwitz zeta function, Rocky Mtn. J. Math. {\bf 2}, 151-157 (1972).}
\bibitem{blago}I. V. Blagouchine, 
{A theorem for the closed–-form evaluation of the first generalized Stieltjes constant at rational arguments, arXiv:1401.3724v2 (2014).}
\bibitem{blago2015}I. V. Blagouchine,
{Expansions of the generalized Euler's constants into the series of polynomials in $\pi^{-2}$ and
into the formal enveloping series with rational coefficients only, arXiv:1501.00740v2.}
\bibitem{briggs}W. E. Briggs, 
{Some constants associated with the Riemann zeta-function, 
Mich. Math. J. {\bf 3}, 117-121 (1955).}
%\bibitem{askey}R. Askey and G. Gasper,
%{J. Anal. Math. {\bf 31}, 48-68 (1977).}
%\bibitem{backlund}R. J. Backlund,
%{\"{U}ber die Nullstellen der Riemannschen Zetafunktion, Acta Math. {\bf 41},
%345-375 (1918).}
%\bibitem{andrews}G. E. Andrews, R. Askey, and R. Roy, Special Functions, Cambridge University
%Press (1999).}
%\bibitem{clark}W. E. Clark and M. E. H. Ismail,
%{Inequalities involving gamma and psi functions,
%http://www.math.usf.edu/\~ clark, to appear in Analysis and Applications.}
%\bibitem{coffey04}M. W. Coffey,
%{Relations and positivity results for derivatives of the Riemann $\xi$
%function, J. Comput. Appl. Math., {\bf 166}, 525-534 (2004).}
\bibitem{coffeyjmaa}M. W. Coffey,
{New results on the Stieltjes constants:  Asymptotic and exact evaluation, 
J. Math. Anal. Appl. {\bf 317}, 603-612 (2006); arXiv:math-ph/0506061.}
\bibitem{coffeystdiffs}M. W. Coffey,
{On representations and differences of Stieltjes coefficients, and other relations,
Rocky Mtn. J. Math. {\bf 41}, 1815-1846 (2011), arXiv/math-ph/0809.3277v2 (2008).}
\bibitem{coffeystv2}M. W. Coffey,
{The Stieltjes constants, their relation to the $\eta_j$ coefficients,
and representation of the Hurwitz zeta function, Analysis {\bf 30}, 383 (2010),
arXiv/math-ph/:0706.0343v2 (2007).}
\bibitem{coffeyseries}M. W. Coffey
{Series representations for the Stieltjes constants, Rocky Mtn. J. Math. {\bf 44}, 443-477
(2014), arXiv/math-ph/0905.1111 (2009).}
%\bibitem{coffeypla}M. W. Coffey,
%{Theta and Riemann xi function representations from harmonic oscillator 
%eigenfunctions, Phys. Lett. A {\bf 362}, 352-356 (2007).}  
\bibitem{coffeydec11}M. W. Coffey,
{Certain logarithmic integrals, including solution of Monthly problem 11629, zeta
values, and expressions for the Stieltjes constants, arXiv:1201.3393 (2012).}
%\bibitem{coffeyaddison}M. W. Coffey,
%{Integral representations of functions and Addison-type series for mathematical constants,
%arXiv:1006.2551 (2010).}
\bibitem{coffeyhypergeo}M. W. Coffey,
{Hypergeometric summation representations of the Stieltjes constants, Analysis {\bf 33}, 
121-142 (2013), arXiv:1106.5148 (2011).}
\bibitem{coffeyrama}M. W. Coffey,
{Functional equations for the Stieltjes constants, to appear in Ramanujan J. (2015), arXiv:1402.3746 (2014).}
%\bibitem{davenport}H. Davenport,
%{Multiplicative Number Theory, Springer Verlag (2000).}
\bibitem{edwards}H. M. Edwards,
{Riemann's Zeta Function, Academic Press, New York (1974).}
%\bibitem{erdelyi}A. Erd\'{e}lyi, W. Magnus, F. Oberhettinger, and G. G. Tricomi,
%{Higher transcendental functions, Vol. I, McGraw-Hill (1953).}
%\bibitem{gould}H. W. Gould,
%{Euler's formula for nth differences of powers, Amer. Math. Monthly
%{\bf 85}, 450-467 (1978).}
\bibitem{grad}I. S. Gradshteyn and I. M. Ryzhik,
{Table of Integrals, Series, and Products, Academic Press, New York (1980).}
%\bibitem{gkp}R. L. Graham, D. E. Knuth, and O. Patashnik,
%{Concrete Mathematics, 2nd ed., Addison Wesley (1994).}  %look up this book 
\bibitem{hansen}E. R. Hansen and M. L. Patrick,
{Some relations and values for the generalized Riemann zeta function,
Math. Comp. {\bf 16}, 265-274 (1962).}
%\bibitem{hardy}G. H. Hardy, 
%{Note on Dr. Vacca's series for $\gamma$, Quart. J. Pure Appl. Math. {\bf 43},
%215-216 (1912).}
%\bibitem{hua}L. K. Hua,
%{Introductiion to Number Theory, Springer Verlag (1982).}
%\bibitem{israilov}M. I. Israilov,
%{On the Laurent Decomposition of Riemann's zeta function,
%Dokl. Akad. Nauk SSSR (Russian) {\bf 12}, 9 (1979); M. I. Israilov,
%Trudy Mat. Inst. Steklova {\bf 158}, 98-104 (1981).}
%\bibitem{ivic}A. Ivi\'{c}, 
%{The Riemann Zeta-Function, Wiley New York (1985).}
%\bibitem{ivic93}A. Ivi\'{c},
%{The Laurent coefficients of certain Dirichlet series,
%Publs. Inst. Math. {\bf 53}, 23-36 (1993).  In Eq. (1.7), $(\log x)^x$ should
%be replaced with $(\log x)^k$ and on p. 25 "comparing with (1.15)" should be 
%replaced with "comparing with (1.5)".}
%\bibitem{jensen}J. L. W. V. Jensen,
%{Sur la fonction $\zeta(s)$ de Riemann, C. R. Acad. Sci. Paris {\bf 104}, 1156-1157
%(1887).}
%\bibitem{kanemitsu}S. Kanemitsu, H. Kumagai, and M. Yoshimoto,
%{Sums involving the Hurwitz zeta function, Ramanujan J. {\bf 5}, 5-19 (2001).}
%\bibitem{karatsuba}A. A. Karatsuba and S. M. Voronin,
%{The Riemann Zeta-Function, Walter de Gruyter, New York (1992).}
%\bibitem{keiper}J. B. Keiper,
%{Power series expansions of Riemann's $\xi$ function, Math. Comp. {\bf 58}, 
%765-773 (1992).}
%\bibitem{knopp}K. Knopp,
%{Theory and Application of Infinite Series, Hafner Publishing Company, New
%York (1951).}
%\bibitem{klusch}D. Klusch,
%{On the Taylor expansion of the Lerch Zeta-function, J. Math. Anal. Appls.
%{\bf 170}, 513-523 (1992).}
\bibitem{kluyver}J. C. Kluyver, 
{On certain series of Mr. Hardy, Quart. J. Pure Appl. Math. {\bf 50}, 185-192
(1927).}
%\bibitem{koepf}W. Koepf and D. Schmersau,
%{Bounded nonvanishing functions and Bateman functions, Complex Variables, 
%{\bf 25}, 237-259 (1994).}
\bibitem{knessl1}C. Knessl and M. W. Coffey,
{An effective asymptotic formula for the Stieltjes constants, Math. Comp. {\bf 80}, 379-386 (2011).}
\bibitem{knessl2}C. Knessl and M. W. Coffey,
{An asymptotic form for the Stieltjes constants $\gamma_k(a)$ and for a sum $S_\gamma(n)$ appearing under the Li criterion, Math. Comp. {\bf 80}, 2197-2217 (2011).}
\bibitem{kreminski}R. Kreminski,
{Newton-Cotes integration for approximating Stieltjes (generalized Euler)
constants, Math. Comp. {\bf 72}, 1379-1397 (2003).}  
%\bibitem{kreminski}R. Kreminski,
%{private communication (2004).}
%\bibitem{lachaud}G. Lachaud,
%{Spectral analysis and the Riemann hypothesis, J. Comp. Appl. Math., {\bf 160},
%175-190 (2003).  On p. 176 of this paper, $\xi(s)$ is missing the factors
%$s(s-1)/2$ and on p. 178 the definition of ${\cal L}$ is misstated.}
%\bibitem{leboeuf}P. Leboeuf and A. G. Monastra,
%{Quantum thermodynamic fluctuations of a chaotic Fermi-gas model, 
%Nucl. Phys. A, {\bf 724}, 69-84 (2003).}
%\bibitem{knuth}D. E. Knuth,
%{The Art of Computer Programming, Vol. 3, Addison Wesley (1973).}  %look up 
%\bibitem{lehmer}D. H. Lehmer,
%{The sum of like powers of the zeros of the Riemann zeta function, Math.
%Comput. {\bf 50}, 265-273 (1988).}
%\bibitem{garunk}A. Laurin\v{c}ikas and R. Garunk\v{s}tis,
%{The Lerch Zeta-function, Kluwer Academic Publishers Dordrecht (2002).}
%\bibitem{lehmer}D. H. Lehmer,
%{The sum of like powers of the zeros of the Riemann zeta function, Math. Comp.
%{\bf 50}, 265-273 (1988).}
%\bibitem{li}X.-J. Li,
%{The positivity of a sequence of numbers and the Riemann hypothesis,
%J. Number Th. {\bf 65}, 325-333 (1997).}
%\bibitem{li2}X.-J. Li,
%{Explicit formulas for Dirichlet and Hecke L-functions, Ill. J. Math., to 
%appear (2004).}
%\bibitem{li2001}X.-J. Li,
%{A formula for the Dedekind $\xi$-function of an imaginary quadratic field, 
%J. Math. Anal. Appl. {\bf 260}, 404-420 (2001).}
%\bibitem{luke}Y. L. Luke,
%{The Special Functions and Their Approximations, Academic Press (1969).}
\bibitem{mitrovic}D. Mitrovi\'{c},
{The signs of some constants associated with the Riemann zeta function,
Mich. Math. J. {\bf 9}, 395-397 (1962).} 
%\bibitem{hlmont}H. L. Montgomery and A. M. Odlyzko,
%{Large deviations of sums of independent random variables, Acta Arith.
%{\bf XLIX}, 427-434 (1988).}
%\bibitem{mati}Yu. V. Matiyasevich,
%{A relationship between certain sums over trivial and nontrivial zeros of the
%Riemann zeta function, Mat. Zametki {\bf 45}, 65-70 (1989).}
%\bibitem{matsuoka}Y. Matsuoka,
%{A note on the relation between generalized Euler constants and the zeros of the
%Riemann zeta function, J. Fac. Educ. Shinshu Univ. {\bf 53}, 81-82 (1985);
%A sequence associated with the zeros of the Riemann zeta function, Tsukuba J.
%Math. {\bf 10}, 249-254 (1986).}
%\bibitem{matsuoka2}Y. Matsuoka,
%{Generalized Euler constants associated with the Riemann zeta function,
%in:  Number Theory and Combinatorics, ed. by J. Akiyama et al., World 
%Scientific, pp. 279-295 (1985).  Concerning just Section 6 here, $t$ should 
%read $T$ on p. 293 (line 4), $\Gamma$ should read $T$ on the same line, and 
%$+$ should read $=$ on the next line.  On the next page, $\gamma_0^3$ should 
%be replaced with $\gamma_0^4$ in Example 1.}
%\bibitem{mestre}J.-F. Mestre,
%{Formules explicites et minorations de conducteurs de vari\'{e}t\'{e}s 
%alg\'{e}briques, Compositio Math. {\bf 58}, 209-232 (1986).}
%\bibitem{cofpla02}M. W. Coffey,
%{Phys. Lett. A {\bf 300}, 367 (2002).}
%\bibitem{neukirch}J. Neukirch,
%{Algebraic Number Theory, Springer Verlag (1999).}
%\bibitem{odlyzko}A. M. Odlyzko,
%{http://www.dtc.umn.edu/\~odlyzko/} %put something else in for this reference?
%\bibitem{amo1989}A. M. Odlyzko,
%{Bounds for discriminants and related estimates for class numbers, regulators 
%and zeros of zeta functions:  A survey of recent results, Sem. Th. No., 
%Bordeaux, {\bf g}, 1-15 (1989).}
%\bibitem{omar}S. Omar,
%{Localization of the first zero of the Dedekind zeta function, Math. Comp.
%{\bf 70}, 1607-1616 (2001).}
%\bibitem{rao}K. S. Rao, G. V. Berghe, and C. Krattenthaler,
%{An entry of Ramanujan on hypergeometric series in his notebooks, preprint
%arXiv:Math.CA/0304317 v1 (2003).}
%\bibitem{other}A. B. Olde Daalhuis and N. M. Temme,
%{Uniform Airy-Type Expansions of Integrals, SIAM J. Math. Anal. {\bf 25}, 
%304-321 (1994).}
%\bibitem{olver}F. W. J. Olver,
%{Asymptotics and Special Functions, Academic Press (1974).}
%\bibitem{planat}M. Planat,
%{1/f noise, the measurement of time and number theory, Fluc. Noise Lett.
%{\bf 1}, R65-R79 (2001).}
%\bibitem{rama}V. Ramaswami,
%{Notes on Riemann's $\zeta$-function, J. London Math. Soc. {\bf 9}, 165-169 
%(1934).}
\bibitem{riemann}B. Riemann,
{\"{U}ber die Anzahl der Primzahlen unter einer gegebenen Gr\"{o}sse, 
Monats. Preuss. Akad. Wiss., 671 (1859-1860).}
%\bibitem{rooney}P. G. Rooney,
%{Further inequalities for generalized Laguerre polynomials, C. R. Math. Rep.
%Acad. Sci. Canada, {\bf 7}, 273-275 (1985).}  %can get without ordering in mail?
\bibitem{stieltjes}T. J. Stieltjes,
{Correspondance d'Hermite et de Stieltjes, Volumes 1 and 2, Gauthier-Villars,
Paris (1905).}
%\bibitem{szego}G. Szeg\"{o},
%{Orthogonal Polynomials, Vol. 23 of AMS Colloquium Publications, AMS (1975).}
%look up this reference:
%\bibitem{temme}N. Temme,
%{Asymptotic estimates for Laguerre polynomials, J. Appl. Math. Phys. {\bf 41},
%114-126 (1990).}  %also to look up:
\bibitem{titch}E. C. Titchmarsh,
{The Theory of the Riemann Zeta-Function, 2nd ed., Oxford University
Press, Oxford (1986).}
%\bibitem{watson}E. T. Whittaker and G. N. Watson,
%{A course of modern analysis, 4th ed., Cambridge Univ. Press, Cambridge (1962).}
%\bibitem{weil}A. Weil,
%{Basic Number Theory, Springer Verlag (1967).}
%\bibitem{weil}A. Weil,
%{Oeuvres Scientifiques--Collected Papers, Vol. II, pp. 48-61, Springer
%(1980).}
\bibitem{wilton2}J. R. Wilton, 
{A note on the coefficients in the expansion of $\zeta(s,x)$ in powers of 
$s-1$, Quart. J. Pure Appl. Math. {\bf 50}, 329-332 (1927).}
%\bibitem{wilton}J. R. Wilton,
%{A proof of Burnside's formula for $\log \Gamma(x+1)$ and certain allied properties 
%of Riemann's $\zeta$-function, Messenger Math. {\bf 52}, 90-93 (1922-1923).}
%\bibitem{wu}H. Wu and D. W. L. Sprung,
%{Riemann zeros and a fractal potential, Phys. Rev. E, {\bf 48}, 2595-2598 
%(1993).}
\bibitem{yue}N.-Y. Zhang and K. S. Williams,
{Some results on the generalized Stieltjes constants, Analysis {\bf 14}, 
147-162 (1994).}  %In addition to a typographical error present in the proof
%of Theorem 5, Remark 5, and Eq. (1.11) of this reference, where $2^{-n}$
%should read $2^{-n-1}$, the first term on the right side of Eq. (6.2) should 
%read $(\log^n a)/a$, and in both Eqs. (6.5) and (6.11), $P_n$ should appear in 
%place of $P_1$.  On p. 148, in Eq. (1.4), $f_n(x)$ should be replaced by
%$f_n'(x)$.  On p. 157, (7.2) should be replaced with 
%(7.1) in the second line of text from the bottom, and on p. 158 (7.2) should 
%be replaced with (1.9) in the third line of text.}  %from the top.}
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphicx}
\usepackage{graphics}
\usepackage{amsthm}
\usepackage{amssymb, amsmath}
\title{Note on the Stieltjes constants: series with Stirling numbers of the first kind}  
\author{Mark W. Coffey\\
Department of Physics\\
Colorado School of Mines\\
Golden, CO  80401\\
mcoffey@mines.edu\\
(Received $\mbox{~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~2015}$)}
\date{April 2, 2015}
\begin{document}
\maketitle
\baselineskip=25 pt
\begin{abstract}

The Stieltjes constants $\gamma_k(a)$ appear as the coefficients in the regular
part of the Laurent expansion of the Hurwitz zeta function $\zeta(s,a)$ about $s=1$.
We generalize the integral and Stirling number series results of \cite{blago2015} for
$\gamma_k(a=1)$.  Along the way, we point out another recent asymptotic development
for $\gamma_k(a)$ which provides convenient and accurate results for even modest values of $k$.


\end{abstract}
 
\vspace{.25cm}
\baselineskip=15pt
\centerline{\bf Key words and phrases}
\medskip 

\noindent

Stieltjes constants, Riemann zeta function, Hurwitz zeta function, Laurent expansion,
digamma function, polygamma function, harmonic numbers 

\vfill
\centerline{\bf 2010 AMS codes}   
11M35, 11M06, 11Y60.  secondary:  05A10

\baselineskip=25pt
\pagebreak
\medskip
\centerline{\bf Introduction and statement of results}
\medskip


The Stieltjes (or generalized Euler) constants $\gamma_k(a)$ appear as 
expansion coefficients in the Laurent series for the Hurwitz zeta function 
$\zeta(s,a)$ about its simple pole at $s=1$ 
\cite{briggs,coffeyjmaa,coffeydec11,kluyver,mitrovic,stieltjes,wilton2},
$$\zeta(s,a)={1 \over {s-1}}+\sum_{n=0}^\infty {{(-1)^n} \over {n!}}\gamma_n(a)(s-1)^n.
\eqno(1.1)$$
These constants are important in analytic number theory and elsewhere, where
they appear in various estimations and as a result of asymptotic analyses, being given
by the limit relation
$$\gamma_k(a)=\lim_{N \to \infty} \left[\sum_{j=1}^N {{\ln^k (j+a)} \over j}-{{\ln^{k+1}(N+a)} \over {k+1}}\right].$$
In particular, $\gamma_0(a)=-\psi(a)$, where $\psi(z)=\Gamma'(z)/\Gamma(z)$ is the
digamma function, with $\Gamma(z)$ the Gamma function.
With $\gamma$ the Euler constant and $\gamma_1=\gamma_1(1)$ and $\gamma_2=\gamma_2(1)$, we
recall the connection with sums of reciprocal powers of the nontrivial zeros $\rho$ of the
Riemann zeta function,
$$\sum_\rho {1 \over \rho^2}=1-{\pi^2 \over 8}+2\gamma_1+\gamma^2, ~~~~
\sum_\rho {1 \over \rho^3}=1-{7 \over 8}\zeta(3)+\gamma^3+3\gamma \gamma_1+{3 \over 2}\gamma_2,$$
such relations following from the Hadamard factorization. 
We recall the connection of differences of Stieltjes constants with logarithmic sums,
$$\gamma_\ell(a)-\gamma_\ell(b)=\sum_{n=0}^\infty \left[{{\ln^\ell(n+a)} \over 
{n+a}}-{{\ln^\ell(n+b)} \over {n+b}}\right].  \eqno(1.2)$$

An effective asymptotic expression for $\gamma_k$ \cite{knessl1} and $\gamma_k(a)$
\cite{knessl2} for $k \gg 1$ has recently been given.  From these expressions, 
which show accurate estimations for even modest values of $k$,
previously known results on sign changes within the sequence of Stieltjes constants follow.
The asymptotic expressions encapsulate both the magnitude $|\gamma_m(a)|$ and 
the changes in sign of the sequence $\{\gamma_m(a)\}$.  Additionally in \cite{knessl2}, an 
elaborate analysis is provided for a certain alternating binomial sum of the Stieltjes constants.

Evaluations of the first and second Stieltjes constants at rational
argument have been given very recently \cite{blago,coffeyrama}.  These decompositions are effectively Fourier series, thus implying many
extensions and applications, and they supplement the relations presented in 
\cite{coffeystdiffs}.  
Besides elaborating on a multiplication formula for the Stieltjes constants, \cite{coffeyrama} provides examples of integrals evaluating in
terms of differences of the first and second of these constants.  In addition, presented
there is a novel method of determining log-log integrals with a certain polynomial 
denominator integrand. 


In this note, we briefly describe how the integral and series representations in \cite{blago2015}
for $\gamma_k(a=1)$ may be readily generalized to $\gamma_k(a)$.  Having the parameter $a$ is
very useful, as even from the $\gamma_0(a)$ case representations then follow for $\ln \Gamma(a)$,
the polygamma functions, and, using integer arguments, harmonic and generalized harmonic numbers.
The key starting point integral representation of these developments is given, so that proofs are
largely omitted.

The Hurwitz zeta function, initially defined by $\zeta(s,a)=\sum_{n=0}^\infty
(n+a)^{-s}$ for $\mbox{Re} ~s>1$, has an analytic continuation to the whole
complex plane \cite{berndt,titch}.   
In the case of $a=1$, $\zeta(s,a)$ reduces to the Riemann zeta function
$\zeta(s)$ \cite{edwards,riemann}.
In this instance, by convention, the Stieltjes constants
$\gamma_k(1)$ are simply denoted $\gamma_k$ \cite{briggs,kluyver,kreminski,mitrovic,yue}.  We recall that
$\gamma_k(a+1)=\gamma_k(a)-(\ln^k a)/a$, and more generally that for $n \geq 1$ an
integer 
$$\gamma_k(a+n)=\gamma_k(a)-\sum_{j=0}^{n-1} {{\ln^k(a+j)} \over {a+j}},$$
as follows from the functional equation $\zeta(s,a+n)=\zeta(s,a)-\sum_{j=0}^{n-1}
(a+j)^{-s}$.  
In fact, an interval of length $1/2$ is sufficient to characterize the $\gamma_k(a)$'s \cite{hansen}.

Unless specified otherwise below, letters $j$, $k$, $\ell$, $m$, $n$, and $r$ denote positive integers. The Euler constant is given by $\gamma=-\psi(1)=\gamma_0(1)$.  The polygamma
functions are denoted $\psi^{(n)}(z)$ and we note that $\psi^{(n)}(z)=(-1)^{n+1}n!\zeta(n+1,z)$
\cite{nbs,grad}.  We let $s(n,k)$ denote the Stirling numbers of the first kind.  We note the
cutoff property $s(n,k)=0$ for $n<k$, consistent with the combinatorial interpretation of
$(-1)^{n+k}s(n,k)$ as the number of permutations of $n$ symbols having $k$ cycles. 


Reference \cite{blago2015} does not seem to be aware of the representation \cite{coffeystv2}
$$\gamma_k(a)={1 \over {2a}}\ln^k a-{{\ln^{k-1}a} \over {k+1}}+{1 \over a}\int_0^\infty
{{(y/a-i)\ln^k(a-iy)+(y/a+i)\ln^k(a+iy)} \over {(1+y^2/a^2)(e^{2\pi y}-1)}}dy. \eqno(1.3)$$
From this integral representation the following results follow.

{\bf Proposition 1}.  Let Re $a>0$.  Then
(a)
$$\gamma_0(a)=-\psi(a)={1 \over {2a}}-\ln a+{1 \over {2\pi ia}}\int_0^1\left[{1 \over {1-{{\ln(1-u)}
\over {2\pi i a}}}}-{1 \over {1+{{\ln(1-u)}\over {2\pi i a}}}}\right]{{du} \over u},$$
(b) 
$$\gamma_0(a)={1 \over {2a}}-\ln a-{1 \over {\pi a}}\sum_{k=0}^\infty {{(2k+1)!(-1)^k} \over 
{(2\pi a)^{2k+1}}}\sum_{n=1}^\infty {{s(n,2k+1)(-1)^n} \over {n! n}},$$
%
and (c)
$$\gamma_0(a)={1 \over {2a}}-\ln a-{1 \over {2\pi^2 a^2}}\sum_{n=0}^\infty {{(2n+1)!(-1)^n} \over 
{(2\pi a)^{2n}}}\zeta(2n+2).$$

{\bf Proposition 2}.  Let Re $a>0$.  Then (a)
$$\gamma_1(a)={{\ln a} \over {2a}}-{1 \over 2}\ln^2a+\ln a\left[-\psi(a)-{1 \over {2a}}+\ln a\right]$$
$$-{1 \over {\pi a}}\sum_{n=1}^\infty {1 \over {nn!}}\sum_{k=0}^{[n/2]}{{(-1)^k} \over {(2\pi a)^{2k+1}}}|s(2k+2,2)| |s(n,2k+1)|,$$
wherein $s(2k+2,2)=(2k+1)!H_{2k+1}$, $H_n=\sum_{k=1}^n 1/k$ being the $n$th harmonic number,
and (b)
$$\gamma_m(a)={{\ln^m a} \over {2a}}-{1 \over {m+1}}\ln^{m+1} a$$
$$+{1 \over {2\pi i}}\int_0^1\left[{{\ln^m\left(a-{{\ln(1-u)} \over {2\pi i}} \right)}\over {a-{{\ln(1-u)}\over {2\pi i}}}}-{{\ln^m\left(a+{{\ln(1-u)} \over {2\pi i}} \right)} \over {a+{{\ln(1-u)}\over {2\pi i}}}}\right]{{du} \over u}.$$

As we discuss, Proposition 1 may be considered the expansion of $\gamma_0(a)$ or of the digamma
function about $|a| \to \infty$ with arg $a<\pi$.  

\medskip
\centerline{\bf Discussion}
\medskip

The proofs of Propositions 1 and 2 follow from (1.3) and the approach of \cite{blago2015}.
Rather than provide such details, we give a complementary illustrating verification of 
Proposition 1.  For this purpose we recall the generating function (e.g., \cite{nbs})
$$\sum_{n=0}^\infty s(n,k){x^n \over {n!}}={1 \over k!}\ln^k(1+x), \eqno(2.1)$$
holding for $|x|<1$ and $|x|=1$ but excluding $x=-1$.   

{\it Proof of Proposition 1}.  By geometric series, the right side of (a) is given by
$${1 \over {2a}}-\ln a+{1 \over {2\pi ia}}\int_0^1\left[{1 \over {1-{{\ln(1-u)}
\over {2\pi i a}}}}-{1 \over {1+{{\ln(1-u)}\over {2\pi i a}}}}\right]{{du} \over u}$$
$$={1 \over {2a}}-\ln a+{2 \over {2\pi ia}}\int_0^1 \sum_{\stackrel{j=0}{j ~\rm{odd}}}^\infty {{\ln^j(1-u)} \over {(2\pi i a)^j}}{{du} \over u},$$
the integral being evaluated with a change of variable,
$$\int_0^1 {{\ln^j(1-u)} \over u}du=(-1)^{j+1}\int_0^\infty {v^j \over {e^v-1}}dv=(-1)^{j+1}
j!\zeta(j+1).$$
Now \cite{grad} (p.\ 943)
$$\psi(z)=\ln z-{1 \over {2a}}-2\int_0^\infty {{t dt} \over {(t^2+z^2)(e^{2\pi t}-1)}}, ~~~~
\mbox{Re} ~z>0.$$
Then again by the use of geometric series,
$$\psi(z)=\ln z-{1 \over {2a}}-{1 \over {2\pi^2 z^2}}\sum_{j=0}^\infty (-1)^j \int_0^\infty
{v^{2j+1} \over {(4\pi^2 z^2)^j}} {{dv} \over {(e^v-1)}}$$
$$=\ln z-{1 \over {2a}}-{1 \over {2\pi^2 z^2}}\sum_{j=0}^\infty {{(-1)^j} \over {(4\pi^2 z^2)^j}}
(2j+1)!\zeta(2j+2).$$
Thus parts (a) and (c) follow.

Part (b) follows from geometric series expansion of the right side of (a) and the use of the
generating function (2.1).  In particular, the term with $s(0,j)=\delta_{0j}$, the Kronecker
symbol, does not contribute. \qed

{\it Remarks}.  By reordering the double sum and using the cutoff property of Stirling numbers 
of the first kind, part (b) may be rewritten as
$$\gamma_0(a)={1 \over {2a}}-\ln a-{1 \over {\pi a}}\sum_{n=1}^\infty {1 \over {n! n}}
\sum_{k=0}^{[n/2]} {{(2k+1)!(-1)^k} \over {(2\pi a)^{2k+1}}} s(n,2k+1)(-1)^n.$$

The Stirling numbers of the first kind may be written with the generalized harmonic
numbers, and the first few are given by $s(n+1,1)=(-1)^n n!$, $s(n+1,2)=(-1)^{n+1}n!H_n$,
$s(n+1,3)=(-1)^n {{n!} \over 2}[H_n^2-H_n^{(2)}]$, and $s(n+1,4)=(-1)^{n+1}{{n!} \over 6}
[H_n^3-3H_nH_n^{(2)}+2H_n^{(3)}]$.    
The generalized harmonic numbers in terms of polygamma function values
$$H_n^{(r)}=\sum_{k=1}^n {1 \over k^r} ={{(-1)^{r-1}} \over {(r-1)!}}\left[\psi^{(r-1)}(n+1)
-\psi^{(r-1)}(1)\right]={{(-1)^{r-1}} \over {(r-1)!}}\int_0^1{{(t^n-1)} \over {t-1}}\ln^{r-1}t
~dt$$  
enter the representations of Proposition 2 for the higher Stieltjes constants.



\pagebreak

\end{document}